Refrigerator A costs $600 and costs $35 a year to operate.
Refrigerator B costs $800 and costs $40 a year to operate.
how many years are the total costs of the 2 regrigerators the same?

Answer by checkley71(8403)

They will never be the same. As the original cost of the most expensive model is greater than the cheaper model & the operating cost of the most expensive model is also greater.
600+35x=800+40x
35x-40x=800-600
-5x=200
x=200/-5
x=-40 non answer.